The Glorious Nebula Resort & Spa expects an earnings per share of $6 in the coming years. The company’s officials do not expect growth in their dividend in the future, and the company pays out 75% of the earnings per share. The investor requires a return rate of 19%. Based on this information, the value of the common stock in the Lord's Treasure Hotel is equal to:

Group of answer choices

$21.50

$23.50

$22.50

$23.68
Answer #1

Dear student thankyou for using homeworklib.

Please find the answer below.

Statement showing computation:.

Dividend= earning*dividend payout ratio

Dividend= 6*0.75= $4.50

Required return= 19%

Current price= Expected Dividend/required return

Current price= 4.50/0.19= $23.68

Current price= $23.68